OpenStudy (anonymous):

what is the slope of the line between (-4,4) and (-1,-2)

OpenStudy (anonymous):

using the change in rise divided by the change in run: -4 - (-1) = -3 for the rise and 4 - (-2) = 6 for the run. -3/6, reduced to -1/2

OpenStudy (anonymous):

(y2-y1) / (x2 - x1)

OpenStudy (anonymous):

so the answer would be 2?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Knowing points (-4, 4) and (-1, -2), the slope, m, is given by: m = [(y2) - (y1)]/[(x2) - (x1)] = [(-2) - (4)]/[(-1) - (-4)] = -6/3 = -2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

np

OpenStudy (anonymous):